yp+x^2q^2=2x^2y


1
Expert's answer
2021-03-10T11:39:00-0500
"yp+x^2q^2=2x^2y\\
\frac{yp}{2x^2y}+\frac{x^2q^2}{2x^2y} =\frac{2x^2y}{2x^2y}\\
\frac{p}{2x^2}+\frac{q^2}{2y} =1\\
\frac{q^2}{2y} =1-\frac{p}{2x^2}=a \qquad \text{(say)}\\
\therefore \qquad \frac{q^2}{2y}=a \implies q = \sqrt{2ay}\\
\text{and } \qquad 1-\frac{p}{2x^2}=a \implies p =2x^2(1-a)"

Since:


"dz = \dfrac{\partial z}{\partial x}dx + \dfrac{\partial z}{\partial y}dy"

Then:


"dz = 2x^2(1-a)\;dx + \sqrt{2ay}\;dy\\"

Integrating through:


"\int dz = \int 2x^2(1-a)\;dx + \int \sqrt{2ay}\;dy\\
z = \frac{2x^3(1-a)}{3} + \frac{\sqrt{8ay^3}}{3}+c\\
3z = 2x^3(1-a) + \sqrt{8ay^3} + b\\"

where b=3c